2)the length of a rectangle is 6 feet more than twice the width. If the length is 24feet, what is the width?
Equation-_________
Answer-__________

Answer by addingup(3677)

L= 2W-6 Now substitute for L:
24= 2W-6 Add 6 to both sides
30= 2W Let's turn it around (more elegant)
2W= 30 Now divide both sides by 2
W= 15 This is the width.
Proof: The problem says that L= 2W-6 and tells us that the L= 24. So:
2(15)-6= 24
30-6= 24
24= 24 We have the correct answer